# Bowen and Wilson Image Class Translation Distance

### v1

@@ Abstract
[Abstract] We propose a novel application of image translation networks for image classification and demonstrate its potential as a more interpretable alternative to conventional black box classification networks. We train a network to translate images between possible classes, and then quantify translation distance, i.e. the degree of alteration needed to conform an image to one class or another. These translation distances can then be examined for clusters and trends, and can be fed directly to a simple classifier (e.g. a support vector machine, SVM), providing comparable accuracy compared to a conventional end-to-end convolutional neural network classifier. In addition, visual inspection of translated images can reveal class-specific characteristics and biases in the training sets, such as visual artifacts that are more frequently observed in one class or another. We demonstrate the approach on a toy 2-class scenario, apples versus oranges, and then apply it to two medical imaging tasks: detecting melanoma from photographs of pigmented lesions and classifying 6 cell types in a bone marrow biopsy smear. This novel application of image-to-image networks shows the potential of the technology to go beyond imagining different stylistic changes and to provide greater insight into image classification and medical imaging datasets.
00footnotetext: title footnote…keywords:
CycleGAN, StarGAN, image classification, image translation, interpretable machine learning, dimension reduction1 Introduction
@@
Recent advances in machine learning such as deep convolutional neural networks (CNNs) have brought about revolutionary new capabilities for automated object recognition and classification 1. While the accuracies yielded are impressive, mistakes and misclassifications are still an area of concern, especially in medical applications. CNNs trained on images from one clinic perform poorly on images from another clinic (known as the out-of-domain problem) and trivial changes to an image, e.g., brightness, lighting, or rotation can fool even state-of-the-art classifiers, even when these changes are incorporated as augmentations to the training data 2. Even worse, misclassifications with high confidence scores can mislead even experts into making the wrong call 3. For these reasons, clinicians have been reluctant to make use of AI in routine practice.
@@
One possible solution is to make a better CNN that does not suffer from the out-of-domain problem, a difficult task that has not yet been solved 4. Another solution lies in interpretable/explainable AI (xAI). The idea is if a human can examine a network’s decision-making process they can decide whether to trust the network’s output for a specific case and use that information to make a more informed decision 5, 6. Most of these approaches work with a conventional CNN classifier and analyze hidden layer activations 7. However, even with advances like class activation maps 8, which enable CNNs to highlight regions that contributed to its decision, a CNN still cannot explain what characteristics led to the decision.
Our approach, illustrated in Fig. 1, is enabled by a generative-adversarial 9 method for learning to translate between different two image domains without the need for paired training images 10 or multiple domains 11. Typical applications of image2image are modality transform (I2I-MT) and class transform (I2I-CT). Examples of I2I-MT include style transfer for artistic purposes, such as transforming a photograph into a cartoon-style image 12 and translating between medical imaging modalities, such as rendering a computed tomography (CT) scan from a magnetic resonance image (MRI) for radiation oncology applications 13. In our application, we use the network (in this study, either a CycleGAN or a StarGAN) for I2I-CT and translate each input image into each possible target class. The amount of alteration can be quantified with pixel-wise mean average distance (eg distance), which can be fed to a classifier such as a support vector machine (SVM) 14. Unlike a conventional black-box classifier, a visual inspection of the translated images can draw an observer’s attention to case-specific defining characteristics and potentially expose biases in the training dataset.
A related approach to ours, "Auditing the inference processes of medical-image classifiers by leveraging generative AI and the expertise of physicians" looked to distinguish between melanomas based on patterns from a generative network 15. Their process used an augmented version of a CycleGAN that produced hypothetical images, that a pretrained classifier would then use for predictions. This approach reveals which alterations (e.g. changes in lighting or background skin tone) can change the classifier’s decision, but does not fix any underlying issues with the traditional network. By contrast, our approach explicitly translates an image between possible classes in order to reveal differences between the classes themselves. The amount of alteration to fit each class is then quantified to digest an image into a simple feature vector that can then be fed to a simple classifier, such as logistic regression.
We emphasize that the main focus of this work is not to outperform state-of-the-art classifiers (although the StarGAN experiments offered significant advantages over a conventional end-to-end CNN classifier). Instead, we aim to show how the image2image framework applied to classification tasks can show highlight the network’s interpretation of an image, visualize morphological changes, and uncover learned domain-specific features, without having to analyze hidden layer activations. This in turn can expose both shortcomings of the network architectures and underlying biases in the datasets, as we will show.
We demonstrate the approach first on two class problems with a CycleGAN: a trivial apples versus oranges task, and then the difficult melanoma versus benign pigmented lesion task. We then compared the performance to a conventional CNN and evaluated the shortcomings of that network architecture, moving into tasks with more than two classes. We then discuss overall results including potential training set biases and shortcomings of the underlying CNNs revealed by visual inspection of the results.
@@ 2 Methods
We conducted four class transformation (I2I-CT) experiments: 1) 2-class CycleGAN on apples/oranges, 2) 2-class CycleGAN pigmented lesion, 3) 3-class StarGAN on bone marrow cytology, and 4) 6-class StarGAN on bone marrow cytology. For each, the performance of the image-to-image approach was compared with a baseline CNN end-to-end classifier (EfficientNet 16, pre-trained on ImageNet 17). For the 2-class problems, to measure performance while accounting for class imbalance we used the area under the receiver operator curve (AUROC) 18. For the multi-class problems, we looked at traditional overall accuracy and then individual class accuracy in a confusion matrix. The resulting generated images were also visually compared with the source images.
2.1 Classification Datasets
2.1.1 2-class problems
The apples/oranges dataset is a subset of the larger ImageNet dataset 19 compiled by the CycleGAN authors 10, from which we randomly selected a training set of 995 apple and 1019 orange images; our test set had 266 apple and 248 orange images.
The pigmented lesions dataset consisted of dermoscopic photographs from the International Skin Imaging Collaboration (ISIC) 2020 challenge dataset 20. Images were collected from the Hospital Clínic de Barcelona, the Medical University of Vienna, Memorial Sloan Kettering Cancer Center, Melanoma Institute Australia, the University of Queensland, and the University of Athens Medical School. There were 30,928 benign images and 584 malignant images. To mitigate the class imbalance problem 21, we used a random subset of 2,584 images with an 80/20 train/test split (453 malignant and 1614 benign images in the training set and 131 malignant and 386 benign images in the test set). The baseline CNN and CycleGAN networks used the same data splits for performance comparisons.
While initial training used the ISIC data from 2020, out-of-distribution performance was tested with malignant images from the 2016 and 2017 ISIC challenges 22, 23, which is comprised of an entirely separate set of images, and additional benign images from the ISIC 2020 challenge that were excluded from our initial subset 20.
2.1.2 Multi-class
For the multi-class experiments, we used the Bone Marrow Cytology in Hematologic Malignancies Data 24. This had 21 different classes. For the 3-class model, we selected three categories: blasts, erythroblasts, and lymphocytes. These classes were selected because of the large availability of images. The number of training images was set at 5,000 images initially for each class (randomly selected), to eliminate the class imbalance problem during training. This ensured that each class was equally likely to be seen during training. The remaining images of each class were used for testing. Training consisted of 15,000 images, 5,000 of each respective class. 50,610 images were used for testing (Blasts: 6,973, Erythroblasts: 22,395, Lymphocytes: 21,242).
For the 6-class model, we used those three classes and added segmented neutrophil, plasma cell, and promyelocyte. These additional three classes were selected because they met the requirements of having more than 5,000 total images. 5,000 images from each of these classes were added to the images of the previous three classes, and the remaining images added to the test set. (Segmented Neutrophil: 24,434, Plasma Cells: 2629, Promylelocyte: 11,994). To test robustness, we wanted to see the capabilities on a larger dataset compared to the 3-class example.
2.2 Baseline CNN classifier
For each dataset, a baseline CNN was implemented using transfer learning. An EffecientNetB3 16 was pre-trained on ImageNet 17. Training augmentations included vertical and horizontal flip, random brightness and contrast, optical distortion, grid distortion, elastic transform, contrast limited adaptive histogram equalization, cutout, and color jitter, using the Albumentation package 25.
To adjust for class imbalance, class weights were used with the datasets with a traditional 80/20 split. This was calculated as . For the datasets with already an equal number of images for each class during, the CNN used the same dataset for comparison purposes with the StarGAN.
The CNN classifiers were trained for fewer iterations than the StarGAN networks due to a CNNs being prone to overtraining. The network used an early stopping callback to monitor the validation set loss with a patience of 200 epochs. The CNN was trained to be a baseline, achieving results that could be comparable to the StarGAN.
@@ 2.3 Image Translation Networks
The CycleGAN model was a TensorFlow implementation 26 of the original PyTorch CycleGAN 10 for 256x256 images. The only modification we applied was the addition of dropout to the residual layers.
Initially, we used the default values of and . Both the generators and discriminators used the optimizer Adam with a learning rate of and . Models were trained for 60,600 iterations.
@@ 2.3.2 Multi-class: StarGAN network
The StarGAN model was the first version of the StarGAN 27. Multi-domain translations are possible with a single network. Labels are given to the different domains and a single generator and discriminator are used to learn the mappings for each image into every domain.
@@
The StarGAN did not originally use an identity loss or calculate translation distances. We implemented a version of identity loss as seen in the CycleGAN 10 and tuned that to improve translation distances and classification. The Bone Cytology and Pigmented lesion experiments used a . (Note: this is smaller than in the CycleGAN because of the difference between the two codebases in using mean absolute error versus sum absolute error in their loss functions). The model was trained for a minimum of 300,000 iterations for the experiments. Increasing the number of classes and samples during training and testing, required more iterations to achieve comparable image quality. A batch size of 16, an image size of 128, and the rest of the default parameters the same as the original StarGAN.
2.4 Image translation distance metrics
To quantify image changes and generate the translation distance feature set, an input image is fed through all generators, producing a set of class-specific hypothetical images, . Using the input images and the generated images, we calculated translation distance (pixel-wise mean average difference) for each class, i.e. , to quantify how much an image needed to change to fit into each class. We expect images to have small translation distances for classes to which they belong and larger translation distances for classes to which they do not belong.
For two classes, the translation distances were visualized with scatterplots, and a translation ratio was calculated, , as a simple estimate of the probability of an image belonging to class . Finally, simple classifier was trained to separate images based on their translation distance coordinates . Most of the experiments used a linear support vector machine (SVM) 14 (with category weighting when dealing with the the pigmented lesions and full bone marrow cytology datasets), but in the 6-class experiment we compare several other classifiers.
The predictions from the CycleGAN+SVM were used to plot the receiver operating characteristic (ROC) and to calculate AUROC as a measurement of the CycleGAN’s ability to separate images by class. As described later, we sought to maximize this CycleGAN+SVM AUROC by tuning the CycleGAN identity loss and cycle consistency loss .
The predictions from the StarGAN+SVM were used to create confusion matrices to monitor class accuracy, specificity, and sensitivity and tune .
@@ 3.1 Apples and Oranges
The separability of apples and oranges in translation distance space can be seen in the scatterplot (Fig. 2(b)) and translation ratio histogram (Fig. 2(c)). The traditional CNN classified apples and oranges with an AUROC of .9922 and .9825 on the training and test sets, respectively. The CycleGAN+SVM achieved an AUROC of .9926 and .9545 on the test and training sets, respectively. By changing to , the AUROC increased to .9971 and .9705 on the training and test sets. Overall, the traditional CNN only outperforms the CycleGAN+SVM approach by a marginal 1.2%, as measured by the test set AUROC, demonstrating the potential of using image translation networks for classification.
Fig. 3 shows a selection of images produced from the apples (class A) to oranges (class B) CyleGAN model with , . The top row of Fig. 3 shows images of reasonable translation quality that were correctly classified by the CycleGAN+SVM. The middle row shows images that failed to translate morphology in the opposing class but translation distances still led to correct classification. The bottom row shows uncharacteristic source images that led to translation or classification failures. Visual inspection shows features such as fruit color, background color, and texture that were changed or added to the original image to produce hypothetical images for each class.
@@ 3.2 Pigmented Lesions
Classifying melanoma versus benign pigmented lesions is relatively more difficult than apples versus oranges. First, we performed hyperparameter tuning to maximize the separation of classes (with an SVM weight of 3 to adjust for the class imbalance). After evaluation models with different combinations of and , we found that and were the best parameters and that identity loss is necessary for preventing the network from changing an image if it is already in the correct domain. (With this class translation task, cycle consistency loss turned out to prevent any meaningful alterations to the images). The CycleGAN+SVM achieved a .8044 AUROC, compared to .8536 AUROC achieved by the CNN classifier. Both approaches are within the range of AUROC reported on the melanoma versus benign dermoscopy task, both for dermatologist visual inspection, where AUROC ranges from 0.671 28 to 0.79 29, with high levels of sensitivity and specificity (70%s and higher), and for CNNs, where AUROC from state-of-the-art networks ranges from 0.654 to 0.949 4 depending on the dataset.
@@
The translation distance scatterplots (a) and (b) in Fig. 4 show that the separation between the malignant and benign pigmented lesions is not as pronounced as in apples versus oranges (Fig. 2). Three regions in translation distance space can be identified: (1) contains benign lesions that are well-separated from the majority of melanomas; (2) contains melanomas that are well-separated from the majority of benign lesions; and (3) contains overlap between benign and malignant lesions. Upon further examination, testing with a larger number of benign images and malignant images from multiple years of the ISIC database (Fig. 5), we found that region (1) consists mostly of images labeled "Single Image Expert Benign" as opposed to "Histopathology Benign" – in other words, these images were considered unambiguously benign by a dermatologist, not requiring a biopsy to confirm. Furthermore, we find that region (1) tends to exclude melanomas, even with a test set expanded to include images from other years of the ISIC database.
@@
For visual inspection, we first show a selection of images (Fig. 6) that were correctly classified by the CycleGAN+SVM at a decision threshold of 0.5. In these cases, rendered images from the correct class tend to look more like the source image in terms of the pigmented lesion’s color (brown/black), texture, and morphology. However, the generators also tend to alter other characteristics such as background skin color (in these cases, pale versus pink), scalebar markings, and vignetting (darkening around the edges due to the imaging optics or lighting conditions). From this, it would appear that the generators have learned to prefer rendering benign lesions with a pink background skin tone, and melanomas with a pale skin tone, scalebar markings, and vignetting. This suggests the presence of these artifacts may be biased towards one class or another in the training data, and may be a source of error in training conventional classifiers.
@@
Secondly, a selection of images misclassified by the traditional CNN is shown in Fig. 9. In these cases, though the CycleGAN+SVM does not always arrive at the correct classification, the generated images provide insight into artifacts and possible causes for misclassification. For example, the benign images in Fig. 9 show the same tendency of the network to consider benign images as having a more pink background tone and melanomas as having lighter skin tone and vignetting. As discussed above, this may reflect biases in the training dataset (further examples can be seen in the supplementary materials). For the malignant images in Fig. 9 (b), while the traditional CNN gave a low probability of melanoma, the generated malignant images appear visually much closer to the source images input malignant lesions, as is also reflected in the translation distances.
@@
Thirdly, a selection of images misclassified by the CycleGAN+SVM is shown in Fig. 9. In these cases, a visual inspection affords more room for interpretation and making an informed decision than would a black-box classifier. Consistent with earlier observations, it appears that benign lesions (Fig. 9 (a)) with pale skin background, scalebar markings, and vignetting tend to be misclassified as melanomas. In the case of malignant lesions (Fig. 9 (b)), though the SVM incorrectly labeled these benign, the source images are more visually similar to the malignant renderings.
@@
Finally, a selection of generator failures/hallucinations are shown in Fig. 9, in which rendered images in one domain or another show severe deviations from the source image. However, in each of these cases, both the CycleGAN+SVM and visual inspection lead to correct classification. Note again the tendency towards pink skin tone in the rendered benign images and pale skin and vignetting in the rendered malignant images.

### v2

# Image class translation: visual inspection of class-specific hypotheticals and classification based on translation distance

@@ Abstract
Purpose:
A major barrier to the implementation of artificial intelligence for medical applications is the lack of explainability and high confidence for incorrect decisions, specifically with out-of-domain samples. We propose a generalization of image translation networks for image classification and demonstrate their potential as a more interpretable alternative to conventional black-box classifiers.
Approach:
We train an image2image network to translate an input image to class-specific hypotheticals, and then compare these with the input, both visually and quantitatively. Translation distances, i.e., the degree of alteration needed to conform to one class or another, are examined for clusters and trends, and used as simple low-dimensional feature vectors for classification.
Results:
On melanoma/benign dermoscopy images, a translation distance classifier achieved 80% accuracy using only a 2-dimensional feature space (versus 85% for a conventional CNN using a -dimensional feature space). Visual inspection of rendered images revealed dataset biases, such as scalebars, vignetting, and pale background pigmentation in melanomas. Image distributions in translation distance space revealed a natural separation along the lines of dermatologist decision to biopsy, rather than between malignant and benign. On bone marrow cytology images, translation distance classifiers outperformed a conventional CNN in both 3-class (92% accuracy vs 89% for CNN) and 6-class (90% vs 86% for CNN) scenarios.
Conclusions:
This proof-of-concept shows the potential for image2image networks to go beyond artistic/stylistic changes and to expose dataset biases, perform dimension reduction and dataset visualization, and in some cases, potentially outperform conventional end-to-end CNN classifiers.
keywords:
[…]
1 Introduction
@@
Recent advances in machine learning such as deep convolutional neural networks (CNNs) have brought about revolutionary new capabilities for automated object recognition and classification [21]. While the accuracies yielded are impressive, mistakes and misclassifications are still an area of concern, especially in medical applications. CNNs trained on images from one clinic perform poorly on images from another clinic (known as the out-of-domain problem) and trivial changes to an image, e.g., brightness, lighting, or rotation can fool even state-of-the-art classifiers, even when these changes are incorporated as augmentations to the training data [1]. Even worse, misclassifications with high confidence scores can mislead even experts into making the wrong call [40]. For these reasons, clinicians have been reluctant to make use of AI in routine practice.
@@
One possible solution is to make a better CNN that does not suffer from the out-of-domain problem, a difficult task that has not yet been solved [43]. Another solution lies in interpretable/explainable AI (xAI). The idea is if a human can examine a network’s decision-making process they can decide whether to trust the network’s output for a specific case and use that information to make a more informed decision [29, 14, 5]. Most of these approaches work with a conventional CNN classifier and analyze hidden layer activations [37]. However, even with advances like class activation maps [47], which highlight pixels that contributed to a decision, a CNN still cannot explain what high-level characteristics were involved.
Here, we propose to use generative-adversarial [15] image2image [48, 8] networks to translate input images into class-specific hypotheticals (where is the number of classes), and compare these with the input – see Fig. 1. Conventional applications of image2image are modality transform (I2I-MT), e.g. style transfer for artistic purposes [7] or translating between medical imaging modalities [23]. The novelty here is using image2image for class transform (I2I-CT) and analyzing, both visually and quantitatively, the changes needed to fit a given image into the possible target classes.
Unlike a conventional black-box classifier, visual inspection of I2I-CT hypotheticals can draw an observer’s attention to case-specific defining characteristics and dataset trends that can bias conventional classifiers (e.g. biologically-irrelevant artifacts such as scalebars and vignetting, which significantly degrade generalizability of CNN classifiers trained to identify melanoma [42, 34]). Quantitative comparison (e.g., through distance between the input and class-specific hypotheticals) reduces each image to a length- vector, , of translation distances. Each dimension of this simple feature vector represents the amount of alteration needed to fit into a class, and can be used to visualize dataset distributions (through, e.g. scatterplots in translation distance space) or as inputs to a simple downstream classifier (e.g. a linear classifier, fully-connected neural network, or support vector machine (SVM) [28]).
Our approach is related to previous works in generative counterfactuals [11], abnormal-to-normal translation (ANT-GAN) [38], and visual attribution ANT-GAN (VANT-GAN) [49]. Whereas saliency and activation mapping approaches (e.g. Grad-CAM [32]) are limited to highlighting which pixels were most important, these generative methods use image2image approaches either to render higher-level factors influencing a classifier’s decision, or to highlight case-specific defining features of disease. In the counterfactuals approach (an extension of Explanation by Progressive Exaggeration [35]), an image2image network learns to produce realistic variations (counterfactual images) that alter a classifier’s decision. A visual inspection of such counterfactuals then reveals whether the classifier is sensitive to the same type of features as human domain experts. In the ANT-GAN approach, an image2image CycleGAN learns to translate images having potential anomalies (e.g. tumors) to hypothetical healthy counterparts. Subtracting these two (or, in the case of VANT-GAN, directly synthesizing the difference image) reveals abnormal regions. More recent work in this space has been to incorporate advances in stable diffusion image generation and large language models [33]. The key differences between counterfactuals and ANT-GAN can be summed up as follows. In counterfactuals, a classifier is central, and image2image is used to analyze the classifier; in ANT-GAN, image2image is central. In counterfactuals, images are produced that either increase or decrease a classifier’s estimated likelihood of membership in a certain class; in ANT-GAN, an image is translated to a possible target class, where ideally translation to the true class leaves the image unaltered.
In this manuscript, we describe an extension and generalization of ANT-GAN, I2I-CT, where (1) target classes can be more than just abnormal versus normal; (2) generated images are not subtracted from source images, rather, visual inspection is performed in like manner to counterfactuals to support interpretation of classifier results and uncover training dataset biases; and (3) translation distances are quantified and evaluated as interpretable low-dimensional representations for images, used for visualizing distributions of images, or as a simple feature vectors for classification. We first demonstrate I2I-CT on two-class problems using CycleGAN: a trivial apples versus oranges task, and then the difficult melanoma versus benign pigmented lesion task. Though a translation distance classifier in these examples does not outperform a conventional CNN (EfficientNet), we include these results to establish proof of principle and highlight the value of manual inspection of class-hypothetical images and translation distance distributions. Finally, we move into and problems with a StarGAN and demonstrate significant overall improvement in accuracy versus a conventional CNN (although we emphasize that classification accuracy improvement over state-of-the-art end-to-end CNNs is not the primary focus of this work).
@@ 2 Methods
I2I-CT uses an image translation network , (e.g. CycleGAN, StarGAN, etc.), to translate an input image to produce a set of corresponding class-specific hypothetical images , where . These images were then compared visually with the original input image, and quantitatively by image translation distance. For proof-of-concept, we simply used pixel-wise mean average difference (though other metrics such as Structural Similarity Index Metric [3] or convolutional encoder perceptual distances [44] could in principle be used). Together, the translation distances for each class-specific hypothetical reduce an image to a simple -dimensional feature vector that can be used for viewing distributions of images (e.g. through scatterplots) in a dataset or directly for classification. The translation distance feature vector has a straightforward interpretation, requiring no technical expertise: each dimension is proportional to the amount of alteration needed to fit an image into a particular class. In other words, images should have small translation distances (ideally, zero) for classes to which they belong and larger translation distances for classes to which they do not belong.
@@
We conducted four class transformation (I2I-CT) experiments: 1) 2-class CycleGAN on apples/oranges, 2) 2-class CycleGAN pigmented lesions, 3) 3-class StarGAN on bone marrow cytology, and 4) 6-class StarGAN on bone marrow cytology. For each, the performance of the image2image approach was compared with a baseline CNN end-to-end classifier (EfficientNet [39], pre-trained on ImageNet [12]). For the 2-class problems, to measure performance while accounting for class imbalance we used the area under the receiver operator curve (AUROC) [18]. For the multi-class problems, we looked at traditional overall accuracy and then individual class accuracy in a confusion matrix.
@@ 2.1 Image Translation Networks
The CycleGAN model was a TensorFlow implementation [26] of the original PyTorch CycleGAN [48] for 256x256 images. The only modification we applied was the addition of dropout to the residual layers. Both the generators and discriminators used the optimizer Adam with a learning rate of and . Models were trained for 60,600 iterations.
The only hyperparameter tuning performed was on the identity loss , which penalizes in-class alterations, and the cycle consistency loss , which encourages reversible transformations. Initially, we used the default values of and . However, we found cycle consistency loss excessively limits morphological modifications such as are expected between melanomas and benign pigmented lesions. These two parameters were tuned as follows. Translation distance vectors were then fed into a SVM, and results were assessed with receiver operating characteristic (ROC) area under the ROC (AUROC) as a measurement of the CycleGAN’s ability to separate images by class. We then sought to maximize AUROC by tuning and .
@@ 2.1.2 Multi-class: StarGAN network
For scenarios, we used the original version of StarGAN [9], to which we added identity loss. As with the CycleGAN, we tuned by evaluating the models based on the resulting accuracy and the confusion matrices of an SVM on the translation distance vectors as well as visual image quality. The Bone Cytology and Pigmented lesion experiments used a . (Note: this is smaller than in the CycleGAN because of the difference between the two codebases in using mean absolute error versus sum absolute error in their loss functions). The model was trained for a minimum of 300,000 iterations for the experiments. Increasing the number of classes and samples during training and testing required more iterations to achieve comparable image quality. We used a batch size of 16, an image size of 128, and the rest of the default parameters the same as the original StarGAN.
2.2 Baseline CNN classifier
For each dataset, a baseline CNN was implemented using transfer learning. An EfficientNetB3 [39] was pre-trained on ImageNet [12]. Training augmentations included vertical and horizontal flip, random brightness and contrast, optical distortion, grid distortion, elastic transform, contrast limited adaptive histogram equalization, cutout, and color jitter, using the Albumentation package [4]. To adjust for class imbalance, class weights were used with the datasets with a traditional 80/20 split. This was calculated as . For the datasets with already an equal number of images for each class during, the CNN used the same dataset for comparison purposes with the StarGAN. The baseline CNNs were trained using an early stopping callback to monitor the validation set loss with a patience of 200 epochs.
2.3 Classification Datasets
2.3.1 2-class datasets
The apples/oranges dataset is a subset of the larger ImageNet dataset [13] compiled by the CycleGAN authors [48], from which we randomly selected a training set of 995 apple and 1019 orange images; our test set had 266 apple and 248 orange images.
The pigmented lesions dataset consisted of dermoscopic photographs from the International Skin Imaging Collaboration (ISIC) 2020 challenge dataset [31], which contains images collected from the Hospital Clínic de Barcelona, the Medical University of Vienna, Memorial Sloan Kettering Cancer Center, Melanoma Institute Australia, the University of Queensland, and the University of Athens Medical School. There were 30,928 benign images and 584 malignant images. To mitigate the class imbalance problem [20], we used a random subset of 2,584 images with an 80/20 train/test split (453 malignant and 1614 benign images in the training set and 131 malignant and 386 benign images in the test set). The baseline CNN and CycleGAN networks used the same data splits for performance comparisons.
2.3.2 Multi-class datasets
For the multi-class experiments, we used the Bone Marrow Cytology in Hematologic Malignancies Data [25], which represents 21 different cell types. For the 3-class model, we selected three categories: blasts, erythroblasts, and lymphocytes. These classes were selected because of the large availability of images. The number of training images was set at 5,000 randomly-selected images for each class (15,000 total training images), to eliminate the class imbalance problem. The remaining 50,610 images of each class were used for testing (6,973 blasts, 22,395 erythroblasts, and 21,242 lymphocytes).
For the 6-class model, we added segmented neutrophils, plasma cells, and promyelocytes. These were selected because they also met the requirements of having more than 5,000 images each. 5,000 images from each of these classes were added to the images of the previous three classes, and the remaining images added to the test set (24,434 segmented neutrophils, 2,629 plasma cells, and 11,994 promylelocytes).
@@ 3.1 Apples and Oranges
The separability of apples and oranges in translation distance space can be seen in the scatterplot (Fig. 2(b)). The end-to-end CNN classified apples and oranges with an AUROC of .9922 and .9825 on the training and test sets, respectively. By contrast, an SVM on translation distance vectors achieved an AUROC of .9926 and .9545 on the training and test sets, respectively. By changing to , the AUROC increased to .9971 and .9705 on the training and test sets. Overall, the traditional CNN only outperforms the CycleGAN+SVM approach by a marginal 1.2%, as measured by the test set AUROC.
Figure 3 shows a selection of images produced from the apples (class 1) to oranges (class 2) CyleGAN model. The top row of Fig. 3 shows images of reasonable translation quality that were correctly classified by a translation distance SVM. The middle row shows images that failed to translate morphology in the opposing class but a translation distance SVM still obtained the correct classification. The bottom row shows uncharacteristic source images that led to translation or classification failures. Visual inspection shows features such as fruit color, background color, and texture that were changed or added to the original image to produce hypothetical images for each class.
@@ 3.2 Pigmented Lesions
Classifying melanoma versus benign pigmented lesions is relatively more difficult than apples versus oranges. First, we performed hyperparameter tuning to maximize the separation of classes (with an SVM weight of 3 to adjust for the class imbalance). After evaluating models with different combinations of and , we found that and were the best parameters and that identity loss is necessary for preventing the network from changing an image if it is already in the correct domain. Image panels generated with different and are shown in Figs. 4 and 5. Nonzero cycle consistency loss turned out to prevent any meaningful alterations to the images. AUROC for different and is shown in Fig. 6(c). The end-to-end CNN achieved 0.8536 test set AUROC, compared to 0.8044 for the translation distance SVM. Both are slightly higher than dermatologist visual inspection, where AUROC ranges from 0.671 [2] to 0.79 [17]. State-of-the-art end-to-end CNNs report test set AUROCs ranging from 0.654 to 0.949, though performance on external datasets is often dramatically lower [43].
@@
The translation distance scatterplots (a) and (b) in Fig. 6 show that the separation between the malignant and benign pigmented lesions is not as pronounced as in apples versus oranges (Fig. 2). Three regions in translation distance space can be identified: (1) contains benign lesions that are well-separated from the majority of melanomas; (2) contains melanomas that are well-separated from the majority of benign lesions; and (3) contains overlap between benign and malignant lesions. Upon further examination, testing with a larger number of benign images and malignant images from multiple years of the ISIC database (Fig. 7), we found that region (1) consists mostly of images labeled “Single Image Expert Benign” as opposed to “Histopathology Benign” – in other words, these images were considered unambiguously benign by a dermatologist, not requiring a biopsy to confirm. Furthermore, we find that region (1) tends to exclude melanomas, even with a test set expanded to include images from other years of the ISIC database.
@@
For visual inspection, we first show a selection of images (Fig. 8) that were correctly classified by the translation distance SVM at a decision threshold of 0.5. In these cases, rendered images from the correct class tend to look more like the source image in terms of the pigmented lesion’s color (brown/black), texture, and morphology. However, the generators also tend to alter other characteristics such as background skin color (in these cases, pale versus pink), scalebar markings, and vignetting (darkening around the edges due to the imaging optics or lighting conditions). From this, it would appear that the generators have learned to prefer rendering benign lesions with a pink background skin tone, and melanomas with a pale skin tone, scalebar markings, and vignetting. This suggests the presence of these artifacts may be biased towards one class or another in the training data, and may be a source of error in training conventional classifiers.
@@
Secondly, a selection of images misclassified by the baseline CNN is shown in Fig. 9. In these cases, though the translation distance SVM does not always arrive at the correct classification, the generated images provide insight into artifacts and possible causes for misclassification. For example, the benign images in Fig. 9 show the same tendency of the network to consider benign images as having a more pink background tone and melanomas as having lighter skin tone and vignetting. As discussed above, this may reflect biases in the training dataset (further examples can be seen in the supplementary materials). For the malignant images in Fig. 9 (b), while the traditional CNN gave a low probability of melanoma, the generated malignant images appear visually much closer to the source images input malignant lesions, as is also reflected in the translation distances. In these cases, a CNN may lead to a missed cancer diagnosis, but visual inspection of I2I-CT hypotheticals would raise concern.
@@
Thirdly, a selection of images misclassified by the translation distance SVM is shown in Fig. 10. In these cases, a visual inspection affords more room for interpretation and making an informed decision than would a black-box classifier. Consistent with earlier observations, it appears that benign lesions (Fig. 10 (a)) with pale skin background, scalebar markings, and vignetting tend to be misclassified as melanomas. In the case of malignant lesions (Fig. 10 (b)), though the SVM incorrectly labeled these benign, the source images are more visually similar to the malignant renderings.
@@
Finally, a selection of generator failures/hallucinations are shown in Fig. 11, in which rendered images in one domain or another show severe deviations from the source image. However, in each of these cases, both the translation distance SVM and visual inspection lead to correct classification. Note again the tendency towards pink skin tone in the rendered benign images and pale skin and vignetting in the rendered malignant images.